# Modeling sea ice in the marginal ice zone as a dense granular flow with rheology inferred from a discrete element model

### v1

@@ 5.2 Variation in ice thickness and number of floes
Next, we test the effectiveness of the continuous model in capturing the DEM results for different ice thicknesses and floe sizes. In figure 11, we plot results for the DEM and the continuous model for steady states with $H=0.5$ and $H=4\,$\mathrm{m}$$, different ocean velocities $u_{o,max}$, and a global sea ice concentration of $A_{0}=0.8$. The ice thickness $H$ enters the continuous model via the parameter $\epsilon=H/L$ in the momentum equation (4.2a). An increase in $H$ is accompanied by an increase in $\epsilon$, which effectively acts as a decrease in the normalized external forcing. We therefore expect an increase in $H$ to result in a decrease in the shear strain and pressure. As observed in panels (a) and (c) of figure 11, this is the case for the velocity profiles of both the continuous model and the DEM; an increase in $H$ is accompanied by a flattening of the normalized velocity profiles. The continuous model is once again accurate in capturing the averaged velocities of the DEM and the general trends in the variation of the sea ice concentration. The pressure resulting from the continuous model and the DEM also decreases as the ice thickness increases, see panel (f); for the pressure, we find a decent fit between the DEM and the continuous model.
Figure 12: Solutions to the non-dimensional continuous system (4.2) (black lines) compared with results from the DEM (markers). For each pair of panels in the first row, we fix the sea ice thickness $H$ and vary the maximum ocean velocity $u_{o,max}$ (the ocean velocity is plotted in blue in panels (a) and (c)). In panel (e), we plot the pressure $p$ in terms of $n$.
Finally, figure 12 contains a comparison between the DEM and the continuous model for different numbers of floes $n$. In particular, we present results with $n=500$, 2000 and 5000 for $A_{0}=0.8$ and $0.9$ and $u_{o,max}=0.5\,$\mathrm{m}$/$\mathrm{s}$$. An increase in $n$ implies a decrease in the effective viscosity of the model, and we therefore expect a decrease in the ice strength or pressure $p$; in the limit where $n\to\infty$, the rheology becomes purely plastic, this can be seen by taking this limit in (4.2a). In figure 12, one can see that the pressure in continuous model indeed decreases with $n$, but this is not the case with the DEM. In fact, the results of the DEM appear to change little with $n$, especially for $A_{0}=0.9$. Despite this, the results from the DEM do not depart from those of the continuous model excessively.
@@ 6 Comparisons with existing models for sea ice
The continuous model studied in this paper shares features with existing models for sea ice. Here, we examine those similarities and also establish differences with our continuous model. Due to its ubiquity in sea ice modeling, we first consider Hibler’s model in section 6.1, before considering other models in section 6.2.
@@ 6.1 Hibler’s model
The most widely used model for sea ice is Hibler’s model, which was first presented in Hibler III, (1979) and treats sea ice like a viscoplastic material. Under the one dimensional conditions of the steady square ocean patch and the non-dimensionalization in section 4, Hibler’s model reduces to the following form:
@@
In (6.1), the parameter $P^{\ast}$ is an empirical constant, $\delta$ is a regularization parameter, and $e=2$ represents the eccentricity of the elliptical yield curve from which Hibler’s model is derived. That is, Hibler’s model assumes a plastic rheology based on an elliptical yield curve which is then regularized. Inside the plane of principal stresses, the yield curve is set in such a way that the ice only resists compression, not pure extension. This geometrical configuration is motivated by the observation that, in pack ice, deformation occurs through ridging (compression) and the opening of leads (extension); of these two mechanisms, only ridging requires a non-negligible amount of plastic work (Rothrock,, 1975).
@@
The main departure between our model and Hibler’s is the expression for the pressure (6.1b): as $u_{o,max}$ increases, $p$ decreases in Hibler’s model, unlike our continuous model where $p$ is independent of $u_{o,max}$. This makes it impossible for Hibler’s model to capture the invariance of the non-dimensional sea ice velocity and pressure with $u_{o,max}$ which we find in most of the DEM results in section 5. This is made clear in figure 13, where we show solutions to Hibler’s model and compare them with the DEM for different values of $A_{0}$ and $u_{o,max}$; an increase in $u_{o,max}$ is accompanied by excessively large changes in the velocity profile. We set $P^{\ast}=5\times 10^{4}\,$\mathrm{N}$$\mathrm{m}$^{-1}$ and $\delta=0.1$ n order to get a good fit with the DEM for some values of $u_{o,max}$. We remark that Hibler’s model is designed for ice that fractures and ridges; in this setting, an increase in the ocean drag weakens the ice through this mechanical deformation. Below, in section 7, we state that future extensions of our continuous model should include the effects of fracturing and ridging, available in SubZero. These future investigations should study the validity of (6.1b) when fracturing and ridging are accounted for.
@@
The choice of an elliptical yield curve in Hibler’s model is motivated by the simplicity of the resulting rheological formulation, yet other possibilities consistent with the requirement of null resistance to pure extension are available, such as the parabolic lens and the teardrop yield curves (Zhang and Rothrock,, 2005; Ringeisen et al.,, 2023) and the Mohr-Coulomb criterion (Ip et al.,, 1991). In fact, as mentioned at the end of section 3, the purely plastic rheology in the $\mu(I)$ formulation considered here (given by (3.2) with $\mu_{1}=0$), can be written as the Mohr-Coulomb rheology presented in Ip et al., (1991) and Gutfraind and Savage, (1997). This becomes clear if we note that $\|\boldsymbol{\mathrm{S}}\|=\mathrm{D}_{1}-\mathrm{D}_{2}$, where $\mathrm{D}_{1}$ and $\mathrm{D}_{2}$ denote the principal components (eigenvalues) of $\boldsymbol{\mathrm{D}}$, and we compare the plastic component in (3.2) with expression (6) in Gutfraind and Savage, (1997). A significant difference between the two expressions is that in (6) from Gutfraind and Savage, (1997) the viscosity is capped to a maximum value in order to avoid the inevitable blow-up that occurs with a purely plastic rheology; this is another form of regularization. This difference of plastic yield curves between Hibler’s model and the purely plastic version of our model is another point of departure between both models in a two dimensional setting.
@@ 6.2 Other models
In section 3, we explain that the viscous component in (3.2) can be interpreted as the rheological effects of collisions, which become increasingly important as $I$ increases. As mentioned in that section, a collisional rheology is derived in Shen et al., (1987) which is also of a linearly viscous nature. In fact, this collisional rheology establishes a viscosity which is linearly proportional to $\overline{d}\sqrt{\rho_{i}Hp}$, as in (3.2) (see Herman, (2022) for a clear description of the collisional rheology). For this, the model proposed in Feltham, (2005) comes close to (3.2), since it takes the ice rheology to be the sum of Hibler’s rheology and the collisional rheology from Shen et al., (1987).
To the knowledge of the authors, the only other study using the $\mu(I)$ rheology to model sea ice is due to Herman, (2022). There, the floes are driven by a moving wall (as opposed to an ocean or atmospheric current, as in our case), and the DEM is based on disk shaped floes, with a much more severe polydispersity. Two main differences can be found between the function $\mu(I)$ that we infer (figure 4) and the one found in Herman, (2022). First, although both cases consider a very similar range of $I$ values, the magnitude of the friction $\mu$ differs considerably, although it is of the same order of magnitude. Second, the $\mu(I)$ curve in Herman, (2022) plateaus for $I>10^{-1}$, something that we do not observe. It remains unclear what may cause these differences. Regarding the second point, we justify our use of a linear function for $\mu$, as in (3.1), by remarking that it simplifies the resulting constitutive equation, enabling a more thorough analysis of the model, and resembles the $\mu$ function proposed in Herman, (2022) over a large range of $I$ values.
@@ 7 Conclusions and future work
In this paper, we have presented a novel mathematical model for sea ice in the MIZ based on the $\mu(I)$ rheology. We have inferred the form of this rheology from DEM computations in section 3. With the analysis in section 4, we prove that the steady one dimensional formulation of this model, given by (4.2), is well-posed in the sense that it has a unique solution. The numerical results presented in Section 5 demonstrate that this model is capable of replicating most of the results of the DEM in the context of steady one dimensional problems. The most visible departure from the continuous model occurs for low ocean velocities and sea ice concentrations, with $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}\leq 0.85$; in this case, the DEM results indicate that the shear stress no longer balances the integrated ocean drag, signaling a breakdown of the underlying continuous momentum balance equation. That is, since inertial effects are found to be negligible in the steady states we consider and the extensional stresses cancel out when integrating across the $x$-direction of the domain to perform the averaging of DEM quantities, our basic continuum model, prior to any choice of rheology, establishes a balance between shear stresses and ocean drag. In figure 10, this balance is found to approximately hold for all cases but those of slow ocean currents and low sea ice concentrations.
@@
The lack of validity of the continuous model for slow ocean currents and low concentrations found in section 5 should be explored further, since this is a regime we expect to find in parts of the MIZ (Stewart et al.,, 2019). As explained in section 5, this lack of accuracy is accompanied by the breakdown of the balance between the averaged shear stress and the integrated ocean drag extracted from the DEM. This balance precedes the choice of rheology and therefore indicates that either a fundamentally different continuous model should be used or some assumption necessary for a continuous model to even hold is no longer valid.
@@
Mechanical processes like fracturing, ridging, and welding are fundamental processes in sea ice dynamics (Feltham,, 2008). Therefore, future improvements of the model presented here should consider the inclusion of these effects. Since ridging and fracturing are fundamental processes in the central ice pack (Rothrock,, 1975), our continuous model could also be valid in this area when these effects are included, yielding a unified sea ice model. Since SubZero includes these mechanical interactions between ice floes (Manucharyan and Montemuro,, 2022), it is a promising tool for exploring their macroscopic effects on the rheology. A first step could be to explore the consequences of including floe fracturing. We expect this would set an upper bound on the pressure and shear stress that the ice cover can sustain. This bound will probably be closely related to the fracture criterion used at the ice floe level. Moreover, floe fracturing will create smaller floes that lead to higher degrees of polydispersity.
@@
The comparisons between the continuous model and the DEM in section 5 are confined to the context of the steady periodic ocean problem, which is essentially one dimensional. Any future studies must examine the accuracy of the $\mu(I)$ rheology in modeling sea ice dynamics under unsteady conditions and in a two-dimensional configuration. Barker et al., (2015) discovered the $\mu(I)$ formulation presented in section 2.1 to be mathematically ill-posed in time-dependent problems, and Schaeffer et al., (2019) has proposed modifications that avoid these instabilities while leaving the steady equations unchanged. A future investigation in the context of sea ice should take these studies into account.
@@
The rheology we propose in (3.2) is local in the sense that the viscosity and the pressure at a certain point of the domain only depend on other quantities and their derivatives at that same point. Yet, granular materials create complex contact networks that enable the interaction of grains set far apart. This leads to non-local effects, and extensions of the $\mu(I)$ rheology that include these effects have been proposed in the context of granular flows (Kamrin and Koval,, 2012; Bouzid et al.,, 2013). Such effects will probably also arise in sea ice modeling and should be considered in future extensions of the model we propose here.
@@ A.1 Computing the stress tensor for an ice floe
where $a_{i}$ is the area of floe $i$, $\boldsymbol{f}^{k}_{i,j}$ is the force at the $k$th contact point exerted by floe $j$ on floe $i$, and $\boldsymbol{r}^{k}_{i,j}$ is the vector connecting the center of mass of floe $i$ with the $k$th contact point. Expression (A.1) corresponds with the Love-Weber formula and, in general, additional terms corresponding to dynamics effects must also be accounted for (Nicot et al.,, 2013). However, we find these dynamic terms to be negligible in all of our computations. We also remark that (A.1) differs from its counterpart in (Manucharyan and Montemuro,, 2022, equation (9)) in two points: (1) we divide by the floe area $a_{i}$ rather than its volume $a_{i}H_{i}$ to obtain the right units and account for the fact that we are working with depth-integrated stresses. (2) We avoid forcing $\boldsymbol{\sigma}_{i}$ to be symmetric and simply use the Love-Weber formula. In general, we find that $\sigma_{i,xy}\approx\sigma_{i,yx}$ in all of our DEM computations.
For the convenience of the reader, we now summarize the calculation of contact forces between two colliding floes in SubZero. A complete account is given in (Manucharyan and Montemuro,, 2022). The contact force $\boldsymbol{f}^{k}_{i,j}$ is the sum of its normal and tangential components,
@@ A.2 Spatial averaging of data
To average the DEM data in space, we must grid the square domain. Taking into account the one dimensional nature of the problem, we divide the domain into $N=10$ cells that stretch in the $x$-direction, such that the edges separating these cells are defined along $N+1$ equispaced points $(y_{0},y_{1},...,y_{N})$ in the $y$-direction. Hence, the resulting grid consists in the cells $[0,L]\times[y_{i-1},y_{i}]$ for $i=1,...,N$, as depicted in figure 15. At each time step, we compute the horizontal velocity $u$ and the components of the stress tensor $\boldsymbol{\sigma}$ by spatially averaging the velocities and stresses of the ice floes contained in each region (the manner in which the stress tensor $\sigma$ is computed for each ice floe is explained in above in appendix A.1. In particular, for each cell of the grid, we perform a mass-weighted averaging such that, for ice floes that are only partially contained in the cell, only the mass of the floe inside the cell is considered. More information about the averaging can be found in Manucharyan and Montemuro, (2022). In order to be consistent with (2.15b), for each run we extract a single pressure $p=\frac{1}{2}(\sigma_{xx}+\sigma_{yy})$ by spatially averaging this quantity over the whole domain.

### v2

@@ 5.2 Variation in ice thickness and number of floes
Next, we test the effectiveness of the continuum model in capturing the DEM results for different ice thicknesses and floe sizes. In figure 11, we plot results for the DEM and the continuum model for steady states with $H=0.5$ and $H=4\,$\mathrm{m}$$, different ocean velocities $u_{o,max}$, and a global sea ice concentration of $A_{0}=0.8$. The ice thickness $H$ enters the continuum model via the parameter $\epsilon=H/L$ in the momentum equation (4.2a). An increase in $H$ is accompanied by an increase in $\epsilon$, which effectively acts as a decrease in the normalized external forcing. We therefore expect an increase in $H$ to result in a decrease in the shear strain and pressure. As observed in panels (a) and (c) of figure 11, this is the case for the velocity profiles of both the continuum model and the DEM; an increase in $H$ is accompanied by a flattening of the normalized velocity profiles. The continuum model is once again accurate in capturing the averaged velocities of the DEM and the general trends in the variation of the sea ice concentration. The pressure resulting from the continuum model and the DEM also decreases as the ice thickness increases, see panel (f); for the pressure, we find a decent fit between the DEM and the continuum model.
Figure 12: Solutions to the non-dimensional continuum model (4.2) (lines) compared with results from the DEM (markers). The solutions to the continuum model are plotted in red for $n=500$, blue for $n=2000$ and green for $n=5000$; the tone is set to light if $A_{0}=0.8$ and dark if $A_{0}=0.9$ (see legend in (e)). For each pair of panels in the first row, we fix the sea ice thickness $H$ and vary the maximum ocean velocity $u_{o,max}$ (the ocean velocity is plotted in blue in panels (a) and (c)). In panel (e), we plot the pressure $p$ in terms of $n$.
Finally, figure 12 contains a comparison between the DEM and the continuum model for different numbers of floes $n$. In particular, we present results with $n=500$, 2000 and 5000 for $A_{0}=0.8$ and $0.9$ and $u_{o,max}=0.5\,$\mathrm{m}$/$\mathrm{s}$$. An increase in $n$ implies a decrease in the effective viscosity of the model, and we therefore expect a decrease in the ice strength or pressure $p$; in the limit where $n\to\infty$, the rheology becomes purely plastic, this can be seen by taking this limit in (4.2a). In figure 12, one can see that the pressure in continuum model indeed decreases with $n$, but this is not the case with the DEM. In fact, the results of the DEM appear to change little with $n$, especially for $A_{0}=0.9$. Despite this, the results from the DEM do not depart from those of the continuum model excessively.
@@ 6 Comparisons with existing models for sea ice
The continuum model studied in this paper shares features with existing models for sea ice. Here, we examine those similarities and also establish differences with our continuum model. Due to its ubiquity in sea ice modeling, we first consider Hibler’s model in section 6.1, before considering other models in section 6.2.
@@ 6.1 Hibler’s model
The most widely used model for sea ice is Hibler’s model, which was first presented in Hibler III (1979) and treats sea ice like a viscoplastic material. Under the one dimensional conditions of the steady square ocean patch and the non-dimensionalization in section 4, Hibler’s model reduces to the following form:
@@
In (6.1), the parameter $P^{\ast}$ is an empirical constant, $\delta$ is a regularization parameter, and $e=2$ represents the eccentricity of the elliptical yield curve from which Hibler’s model is derived. That is, Hibler’s model assumes a plastic rheology based on an elliptical yield curve which is then regularized. Inside the plane of principal stresses, the yield curve is set in such a way that the ice only resists compression, not pure extension. This geometrical configuration is motivated by the observation that, in pack ice, deformation occurs through ridging (compression) and the opening of leads (extension); of these two mechanisms, only ridging requires a non-negligible amount of plastic work (Rothrock, 1975).
@@
The main departure between our model and Hibler’s is the expression for the pressure (6.1b): as $u_{o,max}$ increases, $p$ decreases in Hibler’s model, unlike our continuum model where $p$ is independent of $u_{o,max}$. This makes it impossible for Hibler’s model to capture the invariance of the non-dimensional sea ice velocity and pressure with $u_{o,max}$ which we find in most of the DEM results in section 5. This is made clear in figure 13, where we show solutions to Hibler’s model and compare them with the DEM for different values of $A_{0}$ and $u_{o,max}$; an increase in $u_{o,max}$ is accompanied by excessively large changes in the velocity profile. We set $P^{\ast}=5\times 10^{4}\,$\mathrm{N}$$\mathrm{m}$^{-1}$ and $\delta=0.1$ n order to get a good fit with the DEM for some values of $u_{o,max}$. We remark that Hibler’s model is designed for ice that fractures and ridges; in this setting, an increase in the ocean drag weakens the ice through this mechanical deformation. Below, in section 7, we state that future extensions of our continuum model should include the effects of fracturing and ridging, available in SubZero. These future investigations should study the validity of (6.1b) when fracturing and ridging are accounted for.
@@
The choice of an elliptical yield curve in Hibler’s model is motivated by the simplicity of the resulting rheological formulation, yet other possibilities consistent with the requirement of null resistance to pure extension are available, such as the parabolic lens and the teardrop yield curves (Zhang & Rothrock, 2005; Ringeisen et al., 2023) and the Mohr-Coulomb criterion (Ip et al., 1991). In fact, as mentioned at the end of section 3, the purely plastic rheology in the $\mu(I)$ formulation considered here (given by (3.2) with $\mu_{1}=0$), can be written as the Mohr-Coulomb rheology presented in Ip et al. (1991) and Gutfraind & Savage (1997). This becomes clear if we note that $\|\boldsymbol{\mathrm{S}}\|=\mathrm{D}_{1}-\mathrm{D}_{2}$, where $\mathrm{D}_{1}$ and $\mathrm{D}_{2}$ denote the principal components (eigenvalues) of $\boldsymbol{\mathrm{D}}$, and we compare the plastic component in (3.2) with expression (6) in Gutfraind & Savage (1997). A significant difference between the two expressions is that in (6) from Gutfraind & Savage (1997) the viscosity is capped to a maximum value in order to avoid the inevitable blow-up that occurs with a purely plastic rheology; this is another form of regularization. This difference of plastic yield curves between Hibler’s model and the purely plastic version of our model is another point of departure between both models in a two dimensional setting.
@@ 6.2 Other models
In section 3, we explain that the viscous component in (3.2) can be interpreted as the rheological effects of collisions, which become increasingly important as $I$ increases. As mentioned in that section, a collisional rheology is derived in Shen et al. (1987) which is also of a linearly viscous nature. In fact, this collisional rheology establishes a viscosity which is linearly proportional to $\overline{d}\sqrt{\rho_{i}Hp}$, as in (3.2) (see Herman (2022) for a clear description of the collisional rheology). For this, the model proposed in Feltham (2005) comes close to (3.2), since it takes the ice rheology to be the sum of Hibler’s rheology and the collisional rheology from Shen et al. (1987).
To the knowledge of the authors, the only other study using the $\mu(I)$ rheology to model sea ice is due to Herman (2022). There, the floes are driven by a moving wall (as opposed to an ocean or atmospheric current, as in our case), and the DEM is based on disk shaped floes, with a much more severe polydispersity. Two main differences can be found between the function $\mu(I)$ that we infer (figure 4) and the one found in Herman (2022). First, although both cases consider a very similar range of $I$ values, the magnitude of the friction $\mu$ differs considerably, although it is of the same order of magnitude. Second, the $\mu(I)$ curve in Herman (2022) plateaus for $I>10^{-1}$, something that we do not observe. It remains unclear what may cause these differences. Regarding the second point, we justify our use of a linear function for $\mu$, as in (3.1), by remarking that it simplifies the resulting constitutive equation, enabling a more thorough analysis of the model, and resembles the $\mu$ function proposed in Herman (2022) over a large range of $I$ values.
@@ 7 Conclusions and future work
In this paper, we have presented a novel mathematical model for sea ice in the MIZ based on the $\mu(I)$ rheology. We have inferred the form of this rheology in section 3 using data produced with DEM computations. With the analysis in section 4, we prove that the steady one dimensional formulation of this model, given by (4.2), is well-posed in the sense that it has a unique solution. The numerical results presented in section 5 demonstrate that this model is capable of replicating most of the results of the DEM in the context of steady one dimensional problems. The most visible departure from the continuum model occurs for low ocean velocities and sea ice concentrations, with $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}\leq 0.85$; in this case, the DEM results indicate that the shear stress no longer balances the integrated ocean drag, signaling a breakdown of the underlying grid-averaged momentum balance equation. That is, since inertial effects are found to be negligible in the steady states we consider and the extensional stresses cancel out when integrating across the $x$-direction of the domain to perform the averaging of DEM quantities, our basic continuum model, prior to any choice of rheology, establishes a balance between shear stresses and ocean drag. In figure 10, this balance is found to approximately hold for all cases but those of slow ocean currents and low sea ice concentrations.
@@
The lack of validity of the continuum model for slow ocean currents and low concentrations found in section 5 should be explored further, since this is a regime we expect to find in parts of the MIZ (Stewart et al., 2019). As explained in section 5, this lack of accuracy is accompanied by the breakdown of the balance between the averaged shear stress and the integrated ocean drag extracted from the DEM. This balance precedes the choice of rheology and therefore indicates that either a fundamentally different continuum model should be used or some assumption necessary for a continuum model to even hold is no longer valid.
@@
Mechanical processes like fracturing, ridging, and welding are fundamental processes in sea ice dynamics (Feltham, 2008). Therefore, future improvements of the model presented here should consider the inclusion of these effects. Since ridging and fracturing are fundamental processes in the central ice pack (Rothrock, 1975), our continuum model could also be valid in this area when these effects are included, yielding a unified sea ice model. Since SubZero includes these mechanical interactions between ice floes (Manucharyan & Montemuro, 2022), it is a promising tool for exploring their macroscopic effects on the rheology. A first step could be to explore the consequences of including floe fracturing. We expect this would set an upper bound on the pressure and shear stress that the ice cover can sustain. This bound will probably be closely related to the fracture criterion used at the ice floe level. Moreover, floe fracturing will create smaller floes that lead to higher degrees of polydispersity.
@@
The comparisons between the continuum model and the DEM in section 5 are confined to the context of the steady periodic ocean problem, which is essentially one dimensional. Any future studies must examine the accuracy of the $\mu(I)$ rheology in modeling sea ice dynamics under unsteady conditions and in a two-dimensional configuration. Barker et al. (2015) discovered the $\mu(I)$ formulation presented in section 2.1 to be mathematically ill-posed in time-dependent problems, and Schaeffer et al. (2019) has proposed modifications that avoid these instabilities while leaving the steady equations unchanged. A future investigation in the context of sea ice should take these studies into account.
@@
The rheology we propose in (3.2) is local in the sense that the viscosity and the pressure at a certain point of the domain only depend on other quantities and their derivatives at that same point. Yet, granular materials create complex contact networks that enable the interaction of grains set far apart. This leads to non-local effects, and extensions of the $\mu(I)$ rheology that include these effects have been proposed in the context of granular flows (Kamrin & Koval, 2012; Bouzid et al., 2013). Such effects will probably also arise in sea ice modeling and should be considered in future extensions of the model we propose here.
@@ A.1 Computing the stress tensor for an ice floe
where $a_{i}$ is the area of floe $i$, $\boldsymbol{f}^{k}_{i,j}$ is the force at the $k$th contact point exerted by floe $j$ on floe $i$, and $\boldsymbol{r}^{k}_{i,j}$ is the vector connecting the center of mass of floe $i$ with the $k$-th contact point. Expression (A.1) corresponds with the Love-Weber formula and, in general, additional terms corresponding to dynamics effects must also be accounted for (Nicot et al., 2013). However, we find these dynamic terms to be negligible in all of our computations. We also remark that (A.1) differs from its counterpart in (Manucharyan & Montemuro, 2022, equation (9)) in two points: (1) we divide by the floe area $a_{i}$ rather than its volume $a_{i}H_{i}$ to obtain the right units and account for the fact that we are working with depth-integrated stresses. (2) We avoid forcing $\boldsymbol{\sigma}_{i}$ to be symmetric and simply use the Love-Weber formula. In general, we find that $\sigma_{i,xy}\approx\sigma_{i,yx}$ in all of our DEM computations.
For the convenience of the reader, we now summarize the calculation of contact forces between two colliding floes in SubZero. A complete account is given in (Manucharyan & Montemuro, 2022). The contact force $\boldsymbol{f}^{k}_{i,j}$ is the sum of its normal and tangential components,
@@ A.2 Spatial averaging of data
To average the DEM data in space, we must grid the square domain. Taking into account the one dimensional nature of the problem, we divide the domain into $N=10$ cells that stretch in the $x$-direction, such that the edges separating these cells are defined along $N+1$ equispaced points $(y_{0},y_{1},...,y_{N})$ in the $y$-direction. Hence, the resulting grid consists in the cells $[0,L]\times[y_{i-1},y_{i}]$ for $i=1,...,N$, as depicted in figure 15. At each time step, we compute the horizontal velocity $u$ and the components of the stress tensor $\boldsymbol{\sigma}$ by spatially averaging the velocities and stresses of the ice floes contained in each region (the manner in which the stress tensor $\sigma$ is computed for each ice floe is explained in above in appendix A.1. In particular, for each cell of the grid, we perform a mass-weighted averaging such that, for ice floes that are only partially contained in the cell, only the mass of the floe inside the cell is considered. More information about the averaging can be found in Manucharyan & Montemuro (2022). In order to be consistent with (2.15b), for each run we extract a single pressure $p=\frac{1}{2}(\sigma_{xx}+\sigma_{yy})$ by spatially averaging this quantity over the whole domain.